Martin and Oscar share I L of milk. If Martin gives 50 mL of his share to Oscar, they will have the same volume of milk. Find the original volume of milk that Oscar gets.
Answer by ikleyn(52778)

.

            It can be solved formally,  but the answer is clear/self-evident from the description:

            ANSWER.   Martin has  550 mL initially;   Oscar has  450 mL initially.

            The formal solution is below.


From the condition, you have this system of equations


    M + O = 1000 mL   (1)

    M - 50 = O + 50   (2)


Add the equations.  You will get


    2M + O - 50 = 1000 + O + 50.


Simplify


    2M = 1000 + 50 + 50

    2M = 1100

     M = 1100/5 = 550.


Then from equation (1)


     O = 1000 - M = 1000 - 550 = 450.


as it was predicted at the very beginning.

Solved, answered, explained and completed.